OpenStudy (anonymous):

johansen is baking a cake for ms.farells b-day the recipe calls for 2/5 cups of white sugar and 1/4 of brown sugar.how much more white sugar does johansen need than brown sugar johansen is baking a cake for ms.farells b-day the recipe calls for 2/5 cups of white sugar and 1/4 of brown sugar.how much more white sugar does johansen need than brown sugar @Mathematics

OpenStudy (anonymous):

3/20

OpenStudy (anonymous):

2/5-1/4

OpenStudy (anonymous):

can u show me how u did it

OpenStudy (anonymous):

can u show me how u did it

OpenStudy (anonymous):

2/5-1/4

OpenStudy (anonymous):

can u do the steps

OpenStudy (anonymous):

lets say we need x more amount of white sugar(w) than brown sugar(b) x+b=w x+1/4=2/5 x=2/5-1/4

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yea